differentiate the following function with respect to x
(1) sin^-1(x/sqrt(x^2+a^2)
(2)sin^-1(2x^2 – 1),0
(3) sin^-1(1-2x^2)0

Grade:12

2 Answers

Rinkoo Gupta
9 years ago

Y=sin-1v(1-x2)

Dy/dx=1/v1-(1-x2) .{-2x/v(1-x2)}

=-1/v(1-x2)

Y=sin-1x/v(x2+a2)

Dy/dx=[1/v1-{x/v(x2+a2)}2].[{v(x2+a2).1-x.2x/v(x2+a2)}/(x2+a2)]

=a/(x2+a2)

Thanks & Regards

Rinkoo Gupta

AskIITians Faculty

Arun Kumar IIT Delhi
9 years ago
Hi
figure out the part why you're getting |a| and not just a.

Thanks & Regards, Arun Kumar, Btech,IIT Delhi, Askiitians Faculty